OWG meeting 8-Dec-2008

From GlueXWiki
Jump to: navigation, search


  • Announcements
  • Review of minutes from 12-Nov-2008 meeting (26-Nov meeting was canceled)
  • Update on fiber test - Beni
  • Electronics update (125 MHz FADC, 250 MHz FADC, CTP, SD, etc) - Fernando, Chris and co.
  • Update on other trigger and DAQ issues - Dave D, Graham, others


Note different date, time, room

Mon 8-Dec-2008 9:30am CC F225


Next Meeting

7-Jan-2009 1:30pm CC F326

New Action Items from this Meeting

  • Test IRMIS with PS magnet and with existing Hall D equipment


Attendees: Elliott W, Beni Z, Elton S, Fernando B, Elke A, Chris C, Carl T, Simon T, Dave A, Mark I, Dave D.

Fiber Test Stand

Beni pointed us to his Wiki site, which has complete documentation on recent results (See the entries for Fiber Test Stand on the BCAL wiki page.

  • Beni tested PMT's and SiPM's with different fibers.
  • Best results with Kuraray green fibers with SiPM.
  • Used 140 ns gate.
  • Dark box is working, very light tight.
  • Still need to get stepper motor control working.


Fernando and Chris reported on progress on many projects. See Electronics Wiki Page for full status information. Also, Chris will soon get all the minutes from the Electronics Group meetings on the Wiki.

  • GAS ASIC V2 submission Feb-2009, chips in May-2009, preamp in Jul-2009.
  • New SiPM's from SensL and Hamamatsu due to arrive soon.
  • FADC 125 delayed, most problems solved, thermal test results good, fab should be in Dec-2008, testing Mar-2009, boards in Jul-2009.
  • FADC 250 full 12-bit board Nov-2009, need to switch chips and redo front-end design due to heating problems.
  • F1TDC V2 scheduled for Sep-2009, chips ordered, arrive May-2009.
  • Trigger boards coming along.
  • All boards in-house are tested and working.
  • Two CTP's, one to be stuffed.
  • Quotes in for SD boards, will build two.
  • TI working well.
  • Integrated testing of partial chain of many boards in Feb-2009 (FADC250, CTP, TI, SD).
  • Software being developed, eventually will be merged into CODA3.
  • Still need to work out best electronics <--> detector mapping.


Dave A. reported on recent DAQ progress.

  • A number of internal formats and conventions settled recently, documentation coming.
  • Max events/block is 256 (using 8 bits).
  • Multiple block reads required if different detector electronics share the same crate.
  • Work on prototype ROC starting.


Elliott reported that work on testing IRMIS is delayed due to major version upgrade. Authors asked us to wait a few weeks so we can test the new version (IRMIS3). Also, suggestion to test IRMIS by entering info for all modules, crates, etc. we currently own, including their locations.